OpenStudy (countrygirl1431):

Evaluate 5-|-6| -1 1 11

OpenStudy (anonymous):

-1

OpenStudy (cookie_2046):

-1

OpenStudy (countrygirl1431):

Solve |3 - x| = 8 {-11, 11} {-5, 11} {-5, 5}

OpenStudy (whovianchick):

|3 - x| = 8 3-x=8 3-x=-8 -x=5 -x=-11 x=-5 x=11

OpenStudy (welshfella):

the 3 - x insidee the vertical lines ( these mean 'absolute value') can be positive or negative so we can write 3 - x = 8 or 3 - x = -8 solve these 2 equations for x